What would be the change in the temperature of boiling if 55 g of nacl is added to 774 of water


1
Expert's answer
2021-01-11T03:46:34-0500

The boiling point elevation can be calculated using the following Clausius-Clapeyron and Raoult laws:

"\Delta T_b = i·K_b·b_B" ,

where "\Delta T_b" is the difference of the boiling point of the solution and the boiling point of the solvent (water), "i" is the Van't Hoff factor, accounting for the dissociation (for NaCl, which is a strong electrolyte, "i=1.9" ), "K_b" is the ebullioscopic constant that is specific for the solvent (for water, its value equals 0.512 °C·kg/mol) and "b_B" is the molality of the solute (NaCl).


The molality of the number of the moles of the solute per kilogram of the solvent:

"b_{NaCl} = \frac{n_{NaCl}}{m_{water}} = \frac{m_{NaCl}}{m_{water}·M_{NaCl}} = \frac{55}{774·10^{-3}·58.44} = 1.216" mol/kg.

In this equation, the molar mass "M_{NaCl}" is used to calculate the number of the moles from the mass of NaCl and equals 58.44 g/mol. Therefore, the unit of mass of NaCl is kept in g, but the mass of the solvent (water) is converted to kg, dividing by a factor of 1000.


Finally, the boiling point elevation is:

"\Delta T_b = 1.9·0.512·1.216 = 1.18" °C.


Answer: when 55 g of nacl is added to 774 g of water, the boiling point elevation is 1.18 °C, or 1.2 °C.